# Demagnetisation effects in single-domain particles beat the classical susceptibility limit

### v1

@@ Abstract
According to the classical laws of magnetism, the shape of magnetically soft objects limits the effective susceptibility. For example, spherical soft magnets cannot display an effective susceptibility larger than 3. Although true for macroscopic multi-domain magnetic materials, we show that magnetic nanoparticles in a single-domain state do not suffer from this limitation. We find that the differences between demagnetisation factors along principal axes are relevant and can influence susceptibility for single-domain particles, but do not limit the susceptibility as in the case for multi-domain particles. We validate this result experimentally on spherical nanoparticles with varying diameter (8 to 150 nm) and varying volume fraction (0.1 to 47 vol%). In agreement with our predictions, we measure susceptibilities largely above 3, in fact up to more than 250, for single-domain particles. Moreover, contrary to an existing model, we find that the susceptibility of non-interacting single-domain particles in a non-magnetic matrix scales simply linearly with the volume fraction of particles.
Soft magnetic materials, Superparamagnetism, Magnetic single-domain particle, High-frequency magnetics, Inductor core, Magnetic composite material
@@
where $N$ is the demagnetisation factor of the particle along the applied field direction and $\chi_{p}$ is the intrinsic susceptibility of the particle. The effective susceptibility for magnetically soft materials is therefore at most 1/$N$, or 3 in case of magnetic spheres where $N=1/3$. Consequently, large effective susceptibilities can only be achieved with particles elongated and aligned along the applied field direction.
Eq. (1) is derived implicitly for multi-domain magnets. We question whether it holds for single-domain particles: as these are always magnetically saturated their susceptibility may not be limited by demagnetization effects in the same way. Along these lines, Skomski et al. (2010) [1] found unusual demagnetization effects in magnetization processes of nanoscale spheres (due to coherent magnetization rotation), but did not explore it further.
For ensembles of magnetic particles, the demagnetization correction is more complex due to the combined contributions from the shape of the individual particles, their interactions, and the shape of the sample [2, 3, 4, 5, 6, 7, 8].
For systems of spherical (multi-domain) particles, a global demagnetisation factor that depends on particle volume fraction, $f$, and on the demagnetisation factors of both spherical particles $N=1/3$ and sample shape $N_{s}$, has been proposed [2, 3, 4, 5, 7, 8]
@@
While this model may apply to ensembles of multi-domain particles [5], it is uncertain if it is valid in case of single-domain particles [5], especially if Eq. (1) does not hold for single-domain particles. With Eq. (2), the effective ensemble susceptibility will be at most $\approx$ 11.5 (assuming $f=0.74$ for close-packed spheres and $N_{\mathrm{s}}=0$).
Susceptibility measurements on systems of single-domain nanoparticles of Fe [9, 6], FeNi [10, 11, 12, 13], FeCo [14, 15], and ferrites [16, 17] indicate that spherical single-domain particles may not be limited by demagnetisation effects in the same way as multi-domain soft spherical magnets, although this is not discussed in these references.
Designing high-susceptibility materials is of great interest. For example, new magnetic composite materials containing single-domain magnetic particles are being investigated for sensors [18, 19, 20] and for micro-inductor cores in power electronics [21, 11, 16, 17, 10, 12, 9, 14, 15, 6, 22, 23, 24, 25, 26, 27]. For these applications, it is crucial to understand the effective magnetic susceptibility of single-domain magnetic particles and composite materials containing them.
@@
Noticeably, this shows that the effective susceptibility for single-domain particles depends on the difference in demagnetisation factors along the principal axes of the given shape. Thus, the shape anisotropy does not limit the effective susceptibility to $1/N$ as it does for the multi-domain case. In detail, Eq. (9) establishes that for ellipsoidal single-domain particles, the susceptibility will only be limited by $1/(N_{ii}-N_{jj})$, which has no upper limit, and for spherical single-domain particles, where $N_{xx}=N_{yy}=N_{zz}=1/3$, $\chi_{\text{p,eff}}$ is not 3, but diverges. Eq. (9) reveals also that, in clear contrast to the multi-domain case, spherical single-domain particles achieve the highest susceptibility.
For sufficiently small single-domain particles, thermal energy can induce superparamagnetism. The particle magnetization, while it remains single-domain, reverses at time scales shorter than the observation time. The characteristic timescale, $\tau$, for relaxation between easy directions, depends on the anisotropy energy relative to thermal energy. The time-averaged magnetisation and susceptibility can be found as for a paramagnetic ion (see, e.g., [28, 26]).
@@
From above it is clear that the susceptibility is $\epsilon_{\textrm{M}}/3$ for a spherical particle, i.e. $K_{\textrm{sh}}=0$, and therefore, the susceptibility is not limited to 3 by demagnetisation effects neither in the superparamagnetic single-domain case. If we increase the shape anisotropy we observe that demagnetisation results in slightly larger susceptibility for fields along the long axis and slightly lower susceptibility for fields perpendicular to the long axis. Hence, demagnetisation in superparamagnetic particles acts in similar way as for the blocked case: the differences in demagnetisation factors along primary axes in the particle are of importance. However, susceptibility is further enhanced in some cases due to the anisotropy.
@@
To test our predictions that effective particle susceptibilities above $1/N$ are possible for single-domain particles, we investigated a series of samples containing spherical, fcc Co particles with sizes ranging between 8 and 150 nm on porous Al2O3 support. Particle sizes and distributions were determined by Transmission Electron Microscopy (TEM). Magnetic properties were measured at 200 °C under reducing conditions (2.4% H2 in Ar) by vibrating sample magnetometry (VSM), with the exception of the sample with 8 nm Co particles that was measured at room temperature, while under same gas conditions, and its particle susceptibility was 1/$T$-corrected to 200 °C. For all samples, the Co had been fully reduced in their VSM holder at 900 °C prior to the measurements. The Co particles had a saturation magnetisation of 160$\pm$3 Am2/kg. Particle volume was found by use of measured moment, above saturation magnetisation and Co density. Particle susceptibility were then calculated from demagnetisation corrected sample susceptibility and volume fraction of particles.
Figure 1: Susceptibility and coercive field of fcc Co particles of varying size on Al2O3 support, measured at 200 °C after full reduction at 900 °C. Major/Minor loops measured at applied fields up to 1 T/10 mT. 8 nm particles were measured at room-temperature (RT) and $\chi$ was corrected as 1/T to 200 °C (indicated by *).
Figure 1 shows the measured particle susceptibility and coercivity for Co particles vs. size at 200 °C. It is seen that particle susceptibility for the single domain particles (8-50 nm diameter) are above 9, proving that demagnetisation does not limit particle susceptibility to 3. The particle susceptibility for the smallest particles is even above 250 at 200 °C (400 at room-temperature), with a small coercivity, indicating superparamagnetism. The peak in coercive field indicates that that particles with diameters between 30-60 nm are blocked. Co below 80 nm in diameter has been found to be single domain [29]. It is noted that for the larger particles (140 nm diameter) the susceptibility does not drop to 3. Most likely, these particles are pseudo single-domain i.e. only support few domains, and not fully fitting for the multi-domain case described by equation (1). In all, the measurements of Fig. 1 support that particle susceptibilities above 3 are possible for single-domain particles in both the blocked and superparamagnetic state.
Having derived the effect of demagnetisation on single particle susceptibility, we turn to the topic of nanocomposite susceptibility for materials containing single-domain magnetic particles in a non-magnetic matrix. The most straightforward description would be a linear model [30], where the intrinsic susceptibility of the composites, $\chi_{\textrm{nc}}$, is simply the particle volume fraction, $f$, times the particle susceptibility
@@
This model is often used implicitly for ferromagnetic fluids or other dilute particle systems [31]. It seems readily adaptable to describe the susceptibility of a system of isolated, single-domain particles where the particle susceptibility is not limited to ${1/N}$, but the validity of the model has not yet been systematically investigated for denser particle systems. Specifically, in cases where inter-particle interactions prevail, composite susceptibility may deviate from linear.
Demagnetization correction for nanocomposite sample shape should still be performed, even if Eq. (15) is valid. That is Eq. (1) is applied to $\chi_{\textrm{nc}}$ of Eq. (15) with the demagnetisation factor $N=N_{\textrm{s}}$ in the applied field direction as
@@
If the model implied by Eq. (15) (and Eq. (16)) is valid for nanocomposites, then it precludes Eq. (2), which corrects for both particle and sample shapes. Normile et al. [32] have used the demagnetization factor from Eq. (2) to derive the intrinsic susceptibility of 8 nm maghemite particles in dense assemblies. However, they got large discrepancies in intrinsic susceptibilities (5 vs. 15) for similar particles in thin and thick disks configuration and similar volume fraction (50 vs 59 vol%), suggesting that Eq. (2) may not be suitable for nanocomposites.
@@
In order to test if Eq. 15 is applicable to systems of isolated single-domain particles we prepared and investigated a series of nanocomposites containing varying fraction of spherical $\gamma$-Fe2O3 (maghemite) particles of 11$\pm$3 nm diameter fixed in a PVA polymer matrix. The nanocomposite samples were cast as discs with a thickness of 100-400 $\mu$m and diameter of 6 mm. Small-angle neutron scattering (SANS) confirmed near-spherical particle shape and particle size distributions obtained from TEM, as well as absence of aggregation in the polymer matrix. Detailed description and characterisation of the synthesised nanocomposites is presented elsewhere [27].
We measured the nanocomposites by VSM with sample disc planes parallel to the applied field. All nanocomposite samples showed Langevin behaviour typical for superparamagnetic nanoparticles fitting with distributions similar to the TEM results. No detectable hysteresis was found with coercive fields below 4$\pm$8 A/m at 298 K. Particle saturation magnetisation of $M_{\textrm{s}}=303$ kA/m was found from liquid samples for which iron content had been quantified by inductively coupled plasma mass spectrometry (ICP-MS) [33]. Magnetic particle content in the composites was calculated by use of the particle saturation magnetisation, matrix density of 1 g/cm3, and maghemite density of 4.88 g/cm3.
Figure 2 shows resulting nanocomposite susceptibilities obtained from the VSM measurements. Here, $\chi_{\textrm{nc,eff}}$ (open circles) is given as-measured, and $\chi_{\textrm{nc}}$ (solid squares) is demagnetisation corrected for nanocomposite sample shape (disk /flat cylinder), cf. Eq. (16). The used sample shape demagnetisation factors are calculated according to Ref. [34]. The relatively smaller $\chi_{\textrm{nc,eff}}$ of the sample with 44 volis due to its thicker disk size, and, when demagnetisation corrected for the sample shape, the nanocomposite material follows a linear trend for $\chi_{\textrm{nc}}$ in correspondence with Eq. 15. The theoretic prediction (full line) in Fig. 2, is based on Eqs. (14) and (15), log-normally distributed maghemite particles with $M_{\textrm{s}}=303$ kA/m [27] and a mean diameter of $11\pm 4.1$ nm. We note that the sample shape demagnetization corrected susceptibilities $\chi_{\textrm{nc}}$ are in agreement with our predictions, despite a slightly larger log-normal distribution width than derived from TEM (4.1 nm vs 3 nm), but this variation is within the experimental uncertainty of the TEM data.
@@
For all the nanocomposites in Fig. 2, the corresponding intrinsic particle susceptibility $\chi_{\textrm{p}}$ (cf. Eq. (15)) is consistently around 25. First, this confirms that $\chi_{\textrm{p}}$ is not limited to 3 for spherical nanoparticles. Second, it exemplifies that Eq. (2) cannot hold, as such consistency in $\chi_{\textrm{p}}$ is not obtainable by Eq. (2). Moreover, the high susceptibilities expose experimentally that the demagnetisation factor in Eq. (2), which correct for both particle shape and sample shape, is not applicable to nanoparticle systems. The measured nanocomposite susceptibilities of 11.0 - 11.7 (Fig. 2), are too high to be explained by the demagnetization factor in Eq. (2): $\chi_{\textrm{nc,eff}}$ could maximally be 5.6 for 46 volof spherical nanoparticles, if using the demagnetisation factor in Eq. (2). Additionally, if the here measured nanocomposite susceptibilities of 4.7-11.7 for the 17-46 volsamples (Fig. 2) were demagnetisation corrected for both particle shape and sample shape, as in Eqs. (1)-(2), then meaningless negative intrinsic susceptibilities, $\chi_{\textrm{nc}}$, of -5.1 to -9.4, are obtained, equivalent to overskewed hysteresis loops [1].
@@
Overall, this shows that the demagnetisation factor given in Eq. (2), which may apply to composites of multi-domain particles [5], is not applicable to systems of isolated single-domain nanoparticles. $\chi_{\textrm{nc,eff}}$ should be corrected for sample shape, but not spherical particle shape. Hence, in order to derive $\chi_{\textrm{nc}}$ from experimental measurements on nanocomposites containing isolated, single-domain particles, we suggest following procedure: for a specific sample (e.g. a thin disc of nanocomposite), the effective susceptibility is measured in a certain direction and the measurement is demagnetisation corrected for the sample shape only.
@@
In general, influences of inter-particle interactions are a potential concern for the design of dense nanoparticle composites. Interactions have been reported in literature to either increase or decrease particle susceptibility [6, 28, 35, 36]. In our test samples containing up to 46 vol11±4.1plus-or-minus114.111$\pm$ 4.111 ± 4.1 nm $\gamma$-Fe2O3 particles, we do not observe clear signs of inter-particle interactions, cf. Eq. (15) applies. We estimate the dipolar interaction between particle pairs in our system (1-46 vol11$\pm$3 nm $\gamma$-Fe2O3) to be on the scale of 50-500 K, using $\mu_{0}m^{2}/(4\pi r_{\textrm{cc}}^{3}k_{B})$, where $r_{\textrm{cc}}$ is the mean center-center distance of the particles. This estimate does not take into account the fluctuating nature of the particle magnetisation and that one particle has several neighbours which may lead to increased/decreased interaction fields. When comparing to the anisotropy barrier ($KV/k_{B}$) which is in the range of 1200 K, the found interaction energy scale for our system may be insignificant. This is also in agreement with the simulations in [37] for similar $\epsilon_{\textrm{k}}$ and $\epsilon_{\textrm{M}}$.
@@
In conclusion, we have demonstrated that the effective single-domain particle susceptibility is not limited by demagnetisation, neither for superparamagnetic nor for blocked single-domain particles. The difference between demagnetisation factors along the applied field direction and the particle easy axis is of importance, but enters mainly like an extra (shape) anisotropy and for spherical particles the susceptibility is not limited by particle demagnetisation. It is noticeably that near-spherical shape is the most favourable for obtaining a high effective susceptibility in single-domain particles, in contrast to multi-domain cases. We have validated experimentally the improvement over the classical susceptibility limit by two nanocomposite systems: spherical fcc Co particles of different sizes verify that neither superparamagnetic nor blocked single-domain particles have a particle susceptibility limited by demagnetisation (i.e. $\chi_{p}\gg 3$); spherical superparamagnetic $\gamma$-Fe2O3 particles in different volume fractions confirm that the conventional particle demagnetisation limit does not apply, and conclude that the nanocomposite susceptibility depends linearly on particle volume fraction for non-interacting nanoparticles. Consequently, Eq. 2 does not apply to single-domain particles. The presented results are important for the design of new soft magnetic materials.

### v2

# The Effect of Demagnetisation on the Susceptibility of Single-Domain Particles

@@ Abstract
According to the classical laws of magnetism, the shape of magnetically soft objects limits the effective susceptibility. For example, spherical soft magnets cannot display an effective susceptibility larger than 3. Although true for macroscopic multi-domain magnetic materials, we explain why magnetic nanoparticles in a single-domain state do not suffer from this limitation. We show that the differences between demagnetisation factors along principal axes are relevant and can influence susceptibility for single-domain particles, but do not limit the susceptibility as in the case for multi-domain particles. We validate this result experimentally on spherical nanoparticles with varying diameter (8 to 150 nm) and varying volume fraction (0.1 to 47 vol%). In agreement with our predictions, we measure susceptibilities largely above 3, in fact up to more than 250, for single-domain particles. Moreover, contrary to an existing model, we find that the susceptibility of non-interacting single-domain particles in a non-magnetic matrix scales simply linearly with the volume fraction of particles.
preprint: AIP/123-QED
@@
where $N$ is the demagnetisation factor of the particle along the applied field direction and $\chi_{p}$ is the intrinsic susceptibility of the particle. The effective susceptibility for magnetically soft materials is therefore at most 1/$N$, or 3 in case of magnetic spheres where $N=1/3$. Consequently, large effective susceptibilities can only be achieved with particles elongated and aligned along the applied field direction. Eq. (1) is derived implicitly for multi-domain magnets. We question whether it holds for single-domain particles: as these are always magnetically saturated their susceptibility may not be limited by demagnetization effects in the same way. Along these lines, Skomski et al. (2010) Skomski et al. (2010) found unusual demagnetization effects in magnetization processes of nanoscale spheres (due to coherent magnetization rotation), but did not explore it further.
For ensembles of magnetic particles, the demagnetization correction is more complex due to the combined contributions from the shape of the individual particles, their interactions, and the shape of the sample Breit (1923); Bleaney and Hull (1941); Skomski, Hadjipanayis, and Sellmyer (2007); Bjørk and Bahl (2013); Kin, Kura, and Ogawa (2016); Sánchez et al. (2017); McCann et al. (2022). For systems of spherical (multi-domain) particles, a global demagnetisation factor that depends on particle volume fraction, $f$, and on the demagnetisation factors of both spherical particles $N=1/3$ and sample shape $N_{s}$, has been proposed Breit (1923); Bleaney and Hull (1941); Skomski, Hadjipanayis, and Sellmyer (2007); Bjørk and Bahl (2013); Sánchez et al. (2017); McCann et al. (2022)
@@
While this model may apply to ensembles of multi-domain particles Bjørk and Bahl (2013), it is uncertain if it is valid in case of single-domain particles Bjørk and Bahl (2013), especially if Eq. (1) does not hold for single-domain particles. With Eq. (2), the effective ensemble susceptibility will be at most $\approx$ 11.5 (assuming $f=0.74$ for close-packed spheres and $N_{\mathrm{s}}=0$). Susceptibility measurements on systems of single-domain nanoparticles of Fe Kura et al. (2012); Kin, Kura, and Ogawa (2016), FeNi Yatsugi et al. (2019); Lu et al. (2011); Liu et al. (2005); Williams et al. (2019), FeCo Kura et al. (2014); Yang et al. (2018), and ferrites Yun et al. (2014, 2016) indicate that spherical single-domain particles may not be limited by demagnetisation effects in the same way as multi-domain soft spherical magnets, although this is not discussed in these references.
Designing high-susceptibility materials is of great interest. For example, new magnetic composite materials containing single-domain magnetic particles are being investigated for sensors Bedanta et al. (2007); Moretti et al. (2017); Peserico et al. (2018) and for micro-inductor cores in power electronics Sullivan et al. (2013); Lu et al. (2011); Yun et al. (2014, 2016); Yatsugi et al. (2019); Liu et al. (2005); Kura et al. (2012, 2014); Yang et al. (2018); Kin, Kura, and Ogawa (2016); Garnero et al. (2019); Hasegawa et al. (2009); Rowe et al. (2015); Sanusi et al. (2025); Zambach et al. (2024, 2025). For these applications, it is crucial to understand the effective magnetic susceptibility of single-domain magnetic particles and composite materials containing them.
@@
Noticeably, this shows that the effective susceptibility for single-domain particles depends on the difference in demagnetisation factors along the principal axes of the given shape.
Along the same lines as above we find, that the susceptibility perpendicular to the the easy axis of a prolate particle, where the shape anisotropy easy axis is set parallel to the magneto-crystalline uniaxial anisotropy easy axis, is
$\chi_{\textrm{p,eff}}$ $=\frac{\chi_{\textrm{p}}^{\prime}}{1+\chi_{\textrm{p}}^{\prime}\left(N_{ii}-N_{jj}\right)},$ (10)
with the perpendicular Stoner-Wohlfarth susceptibility $\chi_{\textrm{p}}^{\prime}=\frac{\mu_{0}M_{\textrm{s}}^{2}}{2K_{\textrm{u}}}$ and the difference between demag factors along the principal axis of the spheroid $N_{ii}-N_{jj}$.
Thus, the shape anisotropy does not limit the effective susceptibility to $1/N$ as it does for the multi-domain case. In detail, Eqs. (9) and (10) establish that for ellipsoidal single-domain particles, the susceptibility will only be limited by $1/(N_{ii}-N_{jj})$ due to demagnetisation effects, which has no upper limit, and for spherical single-domain particles, where $N_{xx}=N_{yy}=N_{zz}=1/3$, $\chi_{\text{p,eff}}$ is not 3, but diverges towards the particle susceptibility from other anisotropies. Eq. (9) reveals also that, in clear contrast to the multi-domain case, spherical single-domain particles achieve the highest susceptibility.
For sufficiently small single-domain particles, thermal energy can induce superparamagnetism. The particle magnetization, while it remains single-domain, reverses at time scales shorter than the observation time. The characteristic timescale, $\tau$, for relaxation between easy directions, depends on the anisotropy energy relative to thermal energy. The time-averaged magnetisation and susceptibility can be found as for a paramagnetic ion (see, e.g., Elfimova, Ivanov, and Camp (2019); Zambach et al. (2024)).
@@
From above it is clear that the susceptibility is $\epsilon_{\textrm{M}}/3$ for a spherical particle, i.e. $K_{\textrm{sh}}=0$, and therefore, neither in the superparamagnetic single-domain case, is the susceptibility limited to 3 by demagnetisation effects. If we increase the shape anisotropy we observe that demagnetisation results in slightly larger susceptibility for fields along the long axis and slightly lower susceptibility for fields perpendicular to the long axis. Hence, demagnetisation in superparamagnetic particles acts in similar way as for the blocked case: the differences in demagnetisation factors along primary axes in the particle are of importance. However, susceptibility is further enhanced in some cases due to the anisotropy.
@@
To test our predictions that effective particle susceptibilities above $1/N$ are possible for single-domain particles, we investigated a series of samples containing spherical, fcc Co particles with sizes (mean diameters) ranging between 8 and 150 nm on porous Al2O3 support. Particle sizes and distributions were determined by Transmission Electron Microscopy (TEM). Magnetic properties were measured at 200 °C under reducing conditions (2.4% H2 in Ar) by vibrating sample magnetometry (VSM), with the exception of the sample with 8 nm Co particles that was measured at room temperature, while under same gas conditions, and its particle susceptibility was 1/$T$-corrected to 200 °C. For all samples, the Co had been fully reduced in their VSM holder at 900 °C prior to the measurements. The Co particles showed saturation magnetisation of 1448 kA/m. Particle susceptibility were then calculated from demagnetisation corrected sample susceptibility and volume fraction of particles.
Figure 1: Susceptibility and coercive field of fcc Co particles of varying size on Al2O3 support, measured at 200 °C after full reduction at 900 °C. Major/Minor loops measured at applied fields up to 1 T/10 mT. Particle sizes were found by TEM after magnetic characterisation. 8 nm particles were measured at room-temperature (RT) and $\chi$ was corrected as 1/$T$ to 200 °C (indicated by *).
Figure 1 shows the measured particle susceptibility and coercivity for Co particles vs. size at 200 °C. It is seen that particle susceptibility for the single domain particles (8-50 nm diameter) are above 9, proving that demagnetisation does not limit particle susceptibility to 3. The particle susceptibility for the smallest particles is even above 250 at 200 °C (400 at room-temperature), with a small coercivity, indicating superparamagnetism. The peak in coercive field indicates that that particles with diameters between 30-60 nm are blocked. Co below 80 nm in diameter has been found to be single domain Krishnan (2010). It is noted that for the larger particles (140 nm diameter) the susceptibility does not drop to 3. Most likely, these particles are pseudo single-domain i.e. only support few domains, and not fully fitting for the multi-domain case described by equation (1). In all, the measurements of Fig. 1 support that particle susceptibilities above 3 are possible for single-domain particles in both the blocked and superparamagnetic state.
Having derived the effect of demagnetisation on single particle susceptibility, we turn to the topic of nanocomposite susceptibility for materials containing single-domain magnetic particles in a non-magnetic matrix. The most straightforward description would be a linear model Elfimova, Ivanov, and Camp (2019), where the intrinsic susceptibility of the composites, $\chi_{\textrm{nc}}$, is simply the particle volume fraction, $f$, times the particle susceptibility
@@
This model is often used implicitly for ferromagnetic fluids or other dilute particle systems Elfimova, Ivanov, and Camp (2019); Bai et al. (2018); Carrey, Mehdaoui, and Respaud (2011). It seems readily adaptable to describe the susceptibility of a system of isolated, single-domain particles where the particle susceptibility is not limited to ${1/N}$, but the validity of the model has not yet been systematically investigated for denser particle systems. Specifically, in cases where inter-particle interactions prevail, composite susceptibility may deviate from linear.
Demagnetization correction for nanocomposite sample shape should still be performed, even if Eq. (16) are valid. That is Eq. (1) is applied to $\chi_{\textrm{nc}}$ of Eq. (16) with the demagnetisation factor $N=N_{\textrm{s}}$ in the applied field direction as
@@
If the model implied by Eq. (16) (and Eq. (17)) is valid for nanocomposites, then it precludes Eq. (2), which corrects for both particle and sample shapes. Normile et al. Normile et al. (2016) have used the demagnetization factor from Eq. (2) to derive the intrinsic susceptibility of 8 nm maghemite particles in dense assemblies. However, they got large discrepancies in intrinsic susceptibilities (5 vs. 15) for similar particles in thin and thick disk samples and similar volume fraction (50 vs 59 vol%), suggesting that Eq. (2) may not be suitable for nanocomposites.
In order to test if Eq. 16 is applicable to systems of isolated single-domain particles we prepared and investigated a series of nanocomposites containing varying fraction of spherical $\gamma$-Fe2O3 (maghemite) particles of 11$\pm$3 nm diameter fixed in a poly-vinyl alcohol polymer matrix. The nanocomposite samples were cast as discs with a thickness of 100-400 $\mu$m and diameter of 6 mm. Small-angle neutron scattering confirmed near-spherical particle shape and particle size distributions obtained from TEM, as well as absence of aggregation in the polymer matrix. Detailed description and characterisation of the synthesised nanocomposites is presented elsewhere Zambach et al. (2025).
We measured the nanocomposites by VSM with sample disc planes parallel to the applied field. All nanocomposite samples showed Langevin behaviour typical for superparamagnetic nanoparticles fitting with distributions similar to the TEM results. No detectable hysteresis was found with coercive fields below 4$\pm$8 A/m at 298 K. Particle saturation magnetisation of $M_{\textrm{s}}=303$ kA/m was found from liquid samples for which iron content had been quantified by inductively coupled plasma mass spectrometry Costo et al. (2019). Magnetic particle content in the composites was calculated by use of the particle saturation magnetisation, matrix density of 1 g/cm3, and maghemite density of 4.88 g/cm3.
Figure 2 shows resulting nanocomposite susceptibilities obtained from the VSM measurements. Here, $\chi_{\textrm{nc,eff}}$ (open circles) is given as-measured, and $\chi_{\textrm{nc}}$ (solid squares) is demagnetisation corrected for nanocomposite sample shape (disk /flat cylinder), cf. Eq. (17). The used sample shape demagnetisation factors are calculated according to Ref. Beleggia et al. (2005). The relatively smaller $\chi_{\textrm{nc,eff}}$ of the sample with 44 volis due to its thicker disk size, and, when demagnetisation corrected for the sample shape, the nanocomposite material follows a linear trend for $\chi_{\textrm{nc}}$ in correspondence with Eq. 16. The theoretic prediction (full line) in Fig. 2, is based on Eqs. (15) and (16), log-normally distributed maghemite particles with $M_{\textrm{s}}=303$ kA/m Zambach et al. (2025) and a mean diameter of $11\pm 4.1$ nm. We note that the sample shape demagnetization corrected susceptibilities $\chi_{\textrm{nc}}$ are in agreement with our predictions, despite a slightly larger log-normal distribution width than derived from TEM (4.1 nm vs 3 nm), but this variation is within the experimental uncertainty of the TEM data.
@@
For all the nanocomposites in Fig. 2, the corresponding intrinsic particle susceptibility $\chi_{\textrm{p}}$ (cf. Eq. (16)) is consistently around 25. First, this confirms that $\chi_{\textrm{p}}$ is not limited to 3 for spherical nanoparticles. Second, it exemplifies that Eq. (2) cannot hold, as such consistency in $\chi_{\textrm{p}}$ is not obtainable by Eq. (2). Moreover, the high susceptibilities expose experimentally that the demagnetisation factor in Eq. (2), which correct for both particle shape and sample shape, is not applicable to nanoparticle systems. The measured nanocomposite susceptibilities of 11.0 - 11.7 (Fig. 2), are too high to be explained by the demagnetization factor in Eq. (2): $\chi_{\textrm{nc,eff}}$ could maximally be 5.6 for 46 volof spherical nanoparticles, if using the demagnetisation factor in Eq. (2). Additionally, if the here measured nanocomposite susceptibilities of 4.7-11.7 for the 17-46 volsamples (Fig. 2) were demagnetisation corrected for both particle shape and sample shape, as in Eqs. (1)-(2), then meaningless negative intrinsic susceptibilities, $\chi_{\textrm{nc}}$, of -5.1 to -9.4, are obtained, equivalent to overskewed hysteresis loops Skomski et al. (2010).
@@
Overall, this shows that the demagnetisation factor given in Eq. (2), which may apply to composites of multi-domain particles Bjørk and Bahl (2013), is not applicable to systems of isolated single-domain nanoparticles. $\chi_{\textrm{nc,eff}}$ should be corrected for sample shape, but not spherical particle shape. Hence, in order to derive $\chi_{\textrm{nc}}$ from experimental measurements on nanocomposites containing isolated, single-domain particles, we suggest following procedure: for a specific sample (e.g. a thin disc of nanocomposite), the effective susceptibility is measured in a certain direction and the measurement is demagnetisation corrected for the sample shape only.
@@
In general, influences of inter-particle interactions are a potential concern for the design of dense nanoparticle composites. Interactions have been reported in literature to either increase or decrease particle susceptibility Kin, Kura, and Ogawa (2016); Elfimova, Ivanov, and Camp (2019); Chantrell et al. (2001); Usov and Serebryakova (2020). In our test samples containing up to 46 vol11±4.1plus-or-minus114.111$\pm$ 4.111 ± 4.1 nm $\gamma$-Fe2O3 particles, we do not observe clear signs of inter-particle interactions, cf. Eq. (16) applies. We estimate the dipolar interaction between particle pairs in our system (1-46 vol11$\pm$3 nm $\gamma$-Fe2O3) to be on the scale of 7-70×$\times 10^{-22}$ J, using $\mu_{0}m^{2}/(4\pi r_{\textrm{cc}}^{3})$, where $r_{\textrm{cc}}$ is the mean center-center distance of the particles. This estimate does not take into account the fluctuating nature of the particle magnetisation and that one particle has several neighbours which may lead to increased/decreased interaction fields. When comparing to the anisotropy barrier ($KV$) which is in the range of 170×$\times 10^{-22}$ J when using $K\approx 25$ kJ/m3, the found interaction energy scale for our system may be insignificant. This is also in agreement with the simulations in Durhuus et al. (2025) for similar $\epsilon_{\textrm{k}}$ and $\epsilon_{\textrm{M}}$.
@@
In conclusion, we have explained why the effective single-domain particle susceptibility is not limited by demagnetisation, neither for superparamagnetic nor for blocked single-domain particles. The difference between demagnetisation factors along the applied field direction and the particle easy axis is of importance, but enters mainly like an extra (shape) anisotropy and for spherical particles the susceptibility is not affected by particle demagnetisation. It is noticeably that near-spherical shape is the most favourable for obtaining a high effective susceptibility in single-domain particles, in contrast to multi-domain cases. We have validated experimentally the improvement over the classical susceptibility limit by two nanocomposite systems: spherical fcc Co particles of different sizes verify that neither superparamagnetic nor blocked single-domain particles have a particle susceptibility limited by demagnetisation (i.e. $\chi_{p}\gg 3$); spherical superparamagnetic $\gamma$-Fe2O3 particles in different volume fractions confirm that the conventional particle demagnetisation limit does not apply, and conclude that the nanocomposite susceptibility depends linearly on particle volume fraction for non-interacting nanoparticles. Consequently, Eq. 2 does not apply to single-domain particles. The presented results are important for the design of new soft magnetic materials.